Two coins are tossed and the random variable Z gives the number of heads. Find the range space, possible outcomes, and give the table of values.




1
Expert's answer
2022-03-28T17:55:30-0400

"The\,\,range\,\,space: HH,HT,TH,TT\\Possible\,\,outcomes: 0,1,2\\P\left( Z=0 \right) =P\left( TT \right) =\frac{1}{4}\\P\left( Z=1 \right) =P\left( HT \right) +P\left( TH \right) =\frac{1}{4}+\frac{1}{4}=\frac{1}{2}\\P\left( Z=2 \right) =P\left( HH \right) =\frac{1}{4}"